OpenStudy (anonymous):

how would you write f(x)= (x+5)^2 + (x-9)^2 in standard form??

OpenStudy (anonymous):

f(x)= (x+5)^2 + (x-9)^2 f(x)=x^2+10x+25+x^2-18x+81 f(x)=2x^2-8x+106

OpenStudy (anonymous):

could you explain that middle part